# Measurement and feedforward correction of the fast phase noise of lasers

### v1

@@ I Introduction
Over recent decades, quantum simulation and computation have advanced considerably across platforms, including trapped ions [1, 2], superconducting qubits [3], photonic qubits [4], and cold atoms [5]. To ensure accurate outcomes of the machine, high-fidelity coherent control is crucial for each platform [6, 7, 8, 9]. While external factors often introduce decoherence, the control mechanisms themselves can also inadvertently contribute. Specifically, in neutral-atom or trapped-ion systems, where laser-driven dynamics of a valence electron play a central role, employing highly stable lasers is crucial. Focusing on the problem of driving transition of neutral atoms to Rydberg states, the dominant decoherence sources [10, 8, 11] include the Doppler effect, spontaneous emissions via the intermediate state used for the two-photon excitation, and laser intensity and phase noise [12, 13]—the latter being the focal point of this article. Intuitively, the laser phase needs to be stable on a timescale of 0.1 to 1 microsecond during which the electron is manipulated (literally, shaken) by the laser field. A naive, but illustrative, requirement to suppress the error of a laser-driven quantum gate at the $10^{-3}$ level (0.1 %) would be to ask for less than $\phi_{\rm rms}=(2\pi)\times 10^{-3}=6$ mrad of phase fluctuation in 1 microsecond. Assuming white frequency noise, this translates to a power spectral density (PSD) of phase noise $S_{\phi}(f)=\phi_{\rm rms}^{2}/f=4\times 10^{-11}$ rad2/Hz at Fourier frequency $f=1$ MHz, or equivalently a white frequency noise of $S_{\nu}(f)=f^{2}S_{\phi}(f)=40$ Hz2/Hz. In fact, driven systems are more robust than this simple estimate, and recent analysis indicate that, at such level of PSD, error of manipulation are already suppress at the 0.01 % level [14, 11].
@@ II OVERVIEW
Figure 1: Measurement and correction of fast phase noise. (a) Schematic of the phase noise eater. (b) Time traces of the measured frequency noise with (in blue) and without (in pink) the feedforward correction. (c) Frequency noise PSD with and without the correction. The dark blue curve is the detector noise floor. A cancellation of more than 30 dB is achieved at a Fourier frequency of 3 MHz, where the noise is suppressed down to 0.5 Hz2/Hz.
@@
We now describe the results in the time domain shown in Fig. 1(b). We use a second interferometer, placed after the phase noise eater, to detect the frequency noise with (blue traces) and without (pink) activating the feedforward path. The output voltage, left vertical axis, is converted into frequency deviation, right vertical axis, by using the interferometer calibrated DC sensibility $s_{0}=3.1$ V/MHz. The left (right) panel shows the interferometer raw output, observed with an oscilloscope set on a microsecond (sub-microsecond) timescale. Clearly, the noise has been largely suppressed indicating good performance of the technique. The limits of the phase noise eater can also be distinguished from small noisy features remaining on very fast scales of tens of nanoseconds. The rms deviation is decreased from 200 kHz to 60 kHz, dominated by the faster noise.
@@
For a more quantitative description, we investigate the correction in the Fourier domain with a spectrum analyzer (SA), whose output is shown in Fig. 1(c). Similar to the time domain, we convert the PSD measured by the SA into a frequency-noise PSD from the frequency-dependent sensibility of the interferometer (detailed later). The laser frequency noise with (without) the active correction is shown in blue (pink). The detector limit is also plotted in dark blue and obtained when no light is injected in the interferometer. The peak at 50 MHz originates from the null sensitivity of the interferometer. The unstabilized output has a white frequency noise of $S_{\nu}=700$ Hz2/Hz, attributed to the fundamental quantum noise (Schawlow-Townes limit) of this laser source. By activating the feedforward correction, the frequency noise is strongly suppressed in the 1 to 10 MHz region. The peak suppression is obtained at 3 to 4 MHz, where the PSD is decreased down to 0.5 Hz2/Hz, which is $31.5$ dB below the initial noise level. The noise is remains below 10 Hz2/Hz over the entire 1 to 10 MHz region, with the correction decreasing outside this range. The performance of the noise eater fully agrees with its model, presented later. Its peak performance lies in the region of maximal sensitivity to frequency noise of fast Rydberg quantum gates [11], whose sensitivity drops quickly above the Rabi frequency of the laser drive (typically 3 - 5 MHz). At lower Fourier frequencies ($<1$ MHz), usual feed-back stabilization techniques are amply sufficient and it is thus acceptable for the feedforward correction to not be operative in this region.
@@
The rest of this paper gives details on the results that have been briefly summarized above. In Section III, we elaborate on the measurement of the phase/frequency noise with a delayed Mach-Zehnder interferometer (MZI), and present frequency noise measurements from a collection of lasers found in the laboratory. Then, in Section IV, we detail the feedforward scheme, perform a quantitative comparison to its model, and discuss its performance relative to other reported results in the litterature. Finally, in Section V, we shine the phase noise-canceled laser to atoms and observe an improvement on a Ramsey interfometry signal.
@@ III.1 Phase noise characterization
with $\phi(t)$ a fluctuating phase, and $\nu_{0}$ the (average) laser frequency. Alternatively, one can also consider the instantaneous laser frequency deviation $\nu(t)=(1/2\pi){\rm d}\phi/{\rm d}t$. We are interested in the variation of these quantities on timescale of 0.1 to 1 microsecond, or equivalently in the Fourier domain, around Fourier frequencies $f$ of 1 to 10 MHz, where the sensitivity of a Rydberg quantum gate to laser frequency noise is largest [11]. The adequate tool to quantify the noise of $\phi$ and $\nu$ is the PSD: $S_{\phi}$ (in rad2/Hz) for the phase noise, or $S_{\nu}=f^{2}S_{\phi}$ (in Hz2/Hz) for the frequency noise. The PSD represents the variance of the signal contains in a unit frequency band of 1 Hz centered around $f$. Throughout this work, we will rather work with $S_{\nu}$ and also note that we use single-sided PSD (defined for positive Fourier frequencies only). The PSD can be input directly into a quantum gate error model based on linear response formalism [11], or used to reconstruct random realizations of phase noise feeding numerical simulations [12, 14].
@@
with $V_{\rm pp}$ the peak-to-peak voltage of the fringe, and $\tau$ the delay between the two arms of the interferometer. The approximation on the second line is obtained when the MZI is operated at its quadrature point, giving maximal sensibility to frequency noise. For phase fluctuation slower than the delay $\tau$, we can approximate:
@@
where we introduced the (DC) sensibility $s_{0}$. The sensibility can be directly and accurately calibrated by measuring the peak-to-peak voltage $V_{\rm pp}$ of the MZI fringes on an oscilloscope, and extracting the delay $\tau$ from the zeroes of the MZI sensibility on a spectrum analyzer, see following discussion and Fig. 2(c,d). Typically, we reach a sensibility of $\sim 3$ V/MHz depending on the choice of photo-detector (e.g., with/without a transimpendance amplifier).
@@
leading to a frequency-dependent sensibility:
@@
We see from this equation, also shown in Fig. 2(d), that the MZI display zeroes of sensibility at multiple of $1/\tau$, as well as an overall $1/f$-drop beyond the first zero. We typically use delays $\tau<50$ ns, such that our region of interest ($f=1-10$ MHz) is well within the first zero, which is especially important for the feedforward correction. However, for phase noise measurement, we can also extract information beyond $1/\tau$ (at the cost of a decreasing sensitivity). We note that the choice of delay $\tau$ is a compromise between sensibility (increasing with $\tau$) and measurement bandwidth (decreasing with $\tau$).
@@
Figure 2: A delayed MZI as a fast phase/frequency noise discriminator. (a) Schematic and (b) picture of the realization with fiberized components. (c) Photo-diode (PD) output when scanning the stretcher. The peak-to-peak voltage $V_{\rm pp}$ is recorded to calibrate the interferometer sensitivity, and the interferometer is then locked at mid-point. (d) The detector output is measured with a spectrum analyzer to give the PSD of voltage noise $S$ (in V2/Hz) as a function of the Fourier frequency $f$. The curves in the semi-log plot correspond to the noise obtained when measuring an ECDL laser (pink), and the detection noise floor set by the transimpedance amplifier (blue). Inset: MZI sensitivity $s$ to frequency-noise as a function of the Fourier frequency. The expected nulls of sensitivity at multiples of $f=1/\tau$, clearly observed in (d), are used to precisely calibrate the MZI delay $\tau=22.4$ ns. (e) Frequency noise PSD obtained after converting the spectrum analyzer output $S$ into a frequency-noise PSD $S_{\nu}=S/s^{2}$ using the calibrated MZI sensitivity.
@@ III.2 Fiber MZI realization
We assembled a fiberized MZI from individual polarization-maintaining (PM) components, see Fig. 2(b). The light is first split by a 50/50 PM coupler (Thorlabs, PN780R5F2) into a long and short arm. In the long arm, we place a delay line and a slow phase actuator (a home-made fiber stretcher with resonance beyond kHz). The two arms are then recombined on a second 50/50 coupler. We obtain interference fringes at the two outputs of the MZI, as shown in Fig. 2(c), with visibility of typically 90 % which are stable over time. The residual 10 % loss of contrast is caused by imperfections of the components causing a slight imbalance of power and polarization between the two arms, but it does not affect the performance of the MZI as a frequency discriminator (only slightly decreases the sensibility). The MZI acts as a frequency discriminator when the two arms are combined with a phase difference of $\pi/2$, i.e., when the interferometer is at its quadrature point at the middle of the fringe.
To lock the interferometer, we implemented a slow, sub-kHz, feedback loop with an Arduino DUE board monitoring the MZI output and controlling the fiber stretcher.
The two outputs of the last coupler are detected with photo-diodes. A balanced detection, where the two photo-currents from each output are substracted, is preferred as it rejects (in first-order) the intensity noise from the laser while doubling the sensitivity to frequency noise. The photo-current can either be read directly on a spectrum analyzer, which allows the lowest measurement noise floor, or input to a transimpendance amplifier. We used the second approach (with the Thorlabs PDB425A) to drive the feedforward path, at the cost of a slight increase of noise floor (see next section). The amplifier bandwidth of 100 MHz is well beyond the MZI first zero of sensitivity.
@@ III.3 Laser frequency noise analysis
We then read-out the detector output with a spectrum analyzer to obtain the PSD of voltage $S$ (in V2/Hz). The zeros of sensibility of the MZI appears strikingly at multiple of $1/\tau$ MHz , allowing to precisely calibrate the MZI delay $\tau=22.4$ ns. Together with the measured peak-to-peak voltage $V_{\rm pp}$, this completes the characterization of the MZI sensibility $s_{0}=V_{\rm pp}\pi\tau$. For the signals shown in Fig. 2, we have $s_{0}=1.57$ V/MHz. This calibration procedure is simple and accurate, and is repeated for each interferometers or lasers used in this work. With this sensitivity, we can now convert $S$ (V2/Hz) into the PSD of frequency noise $S_{\nu}=R/s^{2}\times S$ (Hz2/Hz). We note that for a careful quantitative measurement, one has to set the spectrum analyzer detection mode to average rms power (and not the default peak mode, or logarithmic average [35]), and to remember that a spectrum analyzer gives single-sided PSD in case one wants to convert it into two-sided PSD.
@@
When using a biased photo-detector for measurement, the electronic noise floor is first fundamentally set by the thermal Johnson noise of the $R=50\,\Omega$ impedance whose PSD is $S=4k_{B}TR=-180$ dBm/Hz ($0.2$ nV/$\sqrt{\rm Hz}$) at room temperature. In practice, this noise source is overwhelmed by the spectrum analyzer electronic noise floor, typically at $S\simeq-160$ dBm/Hz (2 nV/$\sqrt{\rm Hz}$). This converts into a frequency noise floor of $S_{\nu}\simeq 5\times 10^{-3}$ Hz2/Hz. When using the transimpendance amplifier, the noise floor increases to typically $0.1$ Hz2/Hz, as shown as a blue curve in Fig. 2(d).
In the Fourier frequency range of 1 to 10 MHz, we reached this detector electronic noise floor. At lower frequencies, we however observed a large increase of noise, peaking at 30 kHz, caused by the electronic feedback loop stabilizing the interferometer. This is a combination of the large electronic noise of the Arduino DUE and a resonance of the fiber stretcher. This could be suppressed with better electronics. We conclude that this simple setup is perfectly adequate for measuring fast phase noise on the microsecond timescale and faster, but not for longer timescale.

### v2

@@ I Introduction
Over recent decades, quantum simulation and computation have advanced considerably across platforms, including trapped ions [1, 2], superconducting qubits [3], photonic qubits [4], and cold atoms [5]. To ensure accurate outcomes of the machine, high-fidelity coherent control is crucial for each platform [6, 7, 8, 9]. While external factors often introduce decoherence, the control mechanisms themselves can also inadvertently contribute. Specifically, in neutral-atom or trapped-ion systems, where laser-driven dynamics of a valence electron play a central role, employing highly stable lasers is crucial. Focusing on the problem of driving transition of neutral atoms to Rydberg states, the dominant decoherence sources [10, 8, 11] include the Doppler effect, spontaneous emissions via the intermediate state used for the two-photon excitation, and laser intensity and phase noise [12, 13]—the latter being the focal point of this article. Intuitively, the laser phase needs to be stable on a timescale of 0.1 to 1 microsecond during which the electron is manipulated (literally, shaken) by the laser field. A naive, but illustrative, requirement to suppress the error of a laser-driven quantum gate at the $10^{-3}$ level (0.1 %) would be to ask for less than $\phi_{\rm rms}=(2\pi)\times 10^{-3}=6$ mrad of phase fluctuation in 1 microsecond. Assuming white frequency noise, this translates to a power spectral density (PSD) of phase noise $S_{\phi}(f)$approx $\phi_{\rm rms}^{2}/f=4\times 10^{-11}$ rad2/Hz at Fourier frequency $f=1$ MHz, or equivalently a white frequency noise of $S_{\nu}(f)=f^{2}S_{\phi}(f)=40$ Hz2/Hz. In fact, driven systems are more robust than this simple estimate, and recent analysis indicate that, at such level of PSD, error of manipulation are already suppress at the 0.01 % level [14, 11].
@@ II OVERVIEW
Figure 1: Measurement and correction of fast phase noise. (a) Schematic of the phase noise eater. (b) Time traces of the measured frequency noise with (in blue) and without (in pink) the feedforward correction. (c) Frequency noise PSD with and without the correction. The dark blue curve is the detection noise floor. A cancellation of more than 30 dB is achieved at a Fourier frequency of 3 MHz, where the noise is suppressed down to 0.5 Hz2/Hz.
@@
We now describe the results in the time domain shown in Fig. 1(b). We use a second interferometer, placed after the phase noise eater, to detect the frequency noise with (blue traces) and without (pink) activating the feedforward path. The output voltage, left vertical axis, is converted into frequency deviation, right vertical axis, by using the interferometer calibrated DC sensitivity $s_{0}=3.1$ V/MHz. The left (right) panel shows the interferometer raw output, observed with an oscilloscope set on a microsecond (sub-microsecond) timescale. Clearly, the noise has been largely suppressed indicating good performance of the technique. The limits of the phase noise eater can also be distinguished from small noisy features remaining on very fast scales of tens of nanoseconds. The rms deviation is decreased from 200 kHz to 60 kHz, dominated by the faster noise.
@@
For a more quantitative description, we investigate the correction in the Fourier domain with a spectrum analyzer (SA), whose output is shown in Fig. 1(c). Similar to the time domain, we convert the PSD measured by the SA into a frequency-noise PSD from the frequency-dependent sensitivity of the interferometer (detailed later). The laser frequency noise with (without) the active correction is shown in blue (pink). The detection limit is also plotted in dark blue and obtained when no light is injected in the interferometer. The peak at 50 MHz originates from the null sensitivity of the interferometer. The unstabilized output has a white frequency noise of $S_{\nu}=700$ Hz2/Hz, attributed to the fundamental quantum noise of this laser source. By activating the feedforward correction, the frequency noise is strongly suppressed in the 1 to 10 MHz region. The peak suppression is obtained at 3 to 4 MHz, where the PSD is decreased down to 0.5 Hz2/Hz, which is $31.5$ dB below the initial noise level. The noise is remains below 10 Hz2/Hz over the entire 1 to 10 MHz region, with the correction decreasing outside this range. The performance of the noise eater fully agrees with its model, presented later. Its peak performance lies in the region of maximal sensitivity to frequency noise of fast Rydberg quantum gates [11], whose sensitivity drops quickly above the Rabi frequency of the laser drive (typically 3 - 5 MHz). At lower Fourier frequencies ($<1$ MHz), usual feed-back stabilization techniques are amply sufficient and it is thus acceptable for the feedforward correction to not be operative in this region.
@@
The rest of this paper gives details on the results that have been briefly summarized above. In Section III, we elaborate on the measurement of the phase/frequency noise with a delayed Mach-Zehnder interferometer (MZI), and present frequency noise measurements from a collection of lasers found in the laboratory. Then, in Section IV, we detail the feedforward scheme, perform a quantitative comparison to its model, and discuss its performance relative to other reported results in the literature. Finally, in Section V, we shine the phase noise-canceled laser to atoms and observe an improvement on a Ramsey interferometry signal.
@@ III.1 Phase noise characterization
with $\phi(t)$ a fluctuating phase, and $\nu_{0}$ the (average) laser frequency. Alternatively, one can also consider the instantaneous laser frequency deviation $\nu(t)=(1/2\pi){\rm d}\phi/{\rm d}t$. We are interested in the variation of these quantities on timescale of 0.1 to 1 microsecond, or equivalently in the Fourier domain, around Fourier frequencies $f$ of 1 to 10 MHz, where the sensitivity of a Rydberg quantum gate to laser frequency noise is largest [11]. The adequate tool to quantify the noise of $\phi$ and $\nu$ is the PSD: $S_{\phi}$ (in rad2/Hz) for the phase noise, or $S_{\nu}=f^{2}S_{\phi}$ (in Hz2/Hz) for the frequency noise. The PSD represents the variance of the signal contained in a unit frequency band of 1 Hz centered around $f$. Throughout this work, we will rather work with $S_{\nu}$ and also note that we use single-sided PSD (defined for positive Fourier frequencies only). The PSD can be input directly into a quantum gate error model based on linear response formalism [11], or used to reconstruct random realizations of phase noise feeding numerical simulations [12, 14].
@@
with $V_{\rm pp}$ the peak-to-peak voltage of the fringe, and $\tau$ the delay between the two arms of the interferometer. The approximation on the second line is obtained when the MZI is operated at its quadrature point, giving maximal sensitivity to frequency noise. For phase fluctuation slower than the delay $\tau$, we can approximate:
@@
where we introduced the (DC) sensitivity $s_{0}$. The sensitivity can be directly and accurately calibrated by measuring the peak-to-peak voltage $V_{\rm pp}$ of the MZI fringes on an oscilloscope, and extracting the delay $\tau$ from the zeroes of the MZI sensitivity on a spectrum analyzer, see following discussion and Fig. 2(c,d). Typically, we reach a sensitivity of $\sim 3$ V/MHz depending on the choice of photo-detector (e.g., with/without a transimpedance amplifier).
@@
leading to a frequency-dependent sensitivity :
@@
We see from this equation, also shown in Fig. 2(d), that the MZI display zeroes of sensitivity at multiple of $1/\tau$, as well as an overall $1/f$-drop beyond the first zero. We typically use delays $\tau<50$ ns, such that our region of interest ($f=1-10$ MHz) is well within the first zero, which is especially important for the feedforward correction. However, for phase noise measurement, we can also extract information beyond $1/\tau$ (at the cost of a decreasing sensitivity). We note that the choice of delay $\tau$ is a compromise between sensitivity (increasing with $\tau$) and measurement bandwidth (decreasing with $\tau$).
@@
Figure 2: A delayed MZI as a fast phase/frequency noise discriminator. (a) Schematic and (b) picture of the realization with fiberized components. (c) Photo-diode (PD) output when scanning the stretcher. The peak-to-peak voltage $V_{\rm pp}$ is recorded to calibrate the interferometer sensitivity, and the interferometer is then locked at mid-point. (d) The detector output is measured with a spectrum analyzer to give the PSD of voltage noise $S$ (in V2/Hz) as a function of the Fourier frequency $f$. The curves in the semi-log plot correspond to the noise obtained when measuring an ECDL laser (pink), the electronic noise floor set by the transimpedance amplifier (gray, solid line), the photon shot noise (gray, dashed line) and the total detection noise floor (dark blue). Inset: MZI sensitivity $s$ to frequency-noise as a function of the Fourier frequency. The expected nulls of sensitivity at multiples of $f=1/\tau$, clearly observed in (d), are used to precisely calibrate the MZI delay $\tau=22.4$ ns. (e) Frequency noise PSD obtained after converting the spectrum analyzer output $S$ into a frequency-noise PSD $S_{\nu}=S/s^{2}$ using the calibrated MZI sensitivity.
@@ III.2 Fiber MZI realization
We assembled a fiberized MZI from individual polarization-maintaining (PM) components, see Fig. 2(b). The light is first split by a 50/50 PM coupler (Thorlabs, PN780R5F2) into a long and short arm. In the long arm, we place a delay line and a slow phase actuator (a home-made fiber stretcher with resonance beyond kHz). The two arms are then recombined on a second 50/50 coupler. We obtain interference fringes at the two outputs of the MZI, as shown in Fig. 2(c), with visibility of typically 90 % which are stable over time. The residual 10 % loss of contrast is caused by imperfections of the components causing a slight imbalance of power and polarization between the two arms, but it does not affect the performance of the MZI as a frequency discriminator (only slightly decreases the sensitivity ). The MZI acts as a frequency discriminator when the two arms are combined with a phase difference of $\pi/2$, i.e., when the interferometer is at its quadrature point at the middle of the fringe.
To lock the interferometer, we implemented a slow, sub-kHz, feedback loop with an Arduino DUE board monitoring the MZI output and controlling the fiber stretcher. We put a $\rm kHz$ low-pass filter (not pictured in Fig. 2) after the Arduino to prevent its electronic noise from exciting the resonance frequency of the fiber stretcher.
The two outputs of the last coupler are detected with photo-diodes. A balanced detection, where the two photo-currents from each output are subtracted, is preferred as it rejects (in first-order) the intensity noise from the laser while doubling the sensitivity to frequency noise. The photo-current can either be read directly on a spectrum analyzer, which allows the lowest measurement noise floor, or input to a transimpedance amplifier. We used the second approach (with the Thorlabs PDB425A) to drive the feedforward path, at the cost of a slight increase of noise floor (see next section). The amplifier bandwidth of 100 MHz is well beyond the MZI first zero of sensitivity.
@@ III.3 Laser frequency noise analysis
We then read-out the detector output with a spectrum analyzer to obtain the PSD of voltage $S$ (in V2/Hz). The zeros of sensitivity of the MZI appears strikingly at multiple of $1/\tau$ MHz , allowing to precisely calibrate the MZI delay $\tau=22.4$ ns. Together with the measured peak-to-peak voltage $V_{\rm pp}$, this completes the characterization of the MZI sensitivity $s_{0}=V_{\rm pp}\pi\tau$. For the signals shown in Fig. 2, we have $s_{0}=1.57$ V/MHz. This calibration procedure is simple and accurate, and is repeated for each interferometers or lasers used in this work. With this sensitivity, we can now convert $S$ (V2/Hz) into the PSD of frequency noise $S_{\nu}=R/s^{2}\times S$ (Hz2/Hz). We note that for a careful quantitative measurement, one has to set the spectrum analyzer detection mode to average rms power (and not the default peak mode, or logarithmic average [35]), and to remember that a spectrum analyzer gives single-sided PSD in case one wants to convert it into two-sided PSD.
@@
When using a biased photo-detector for measurement, the electronic noise floor is first fundamentally set by the thermal Johnson noise of the $R=50\,\Omega$ impedance whose PSD is $S=4k_{B}TR=-180$ dBm/Hz ($0.2$ nV/$\sqrt{\rm Hz}$) at room temperature. In practice, this noise source is overwhelmed by the spectrum analyzer electronic noise floor, typically at $S\simeq-160$ dBm/Hz (2 nV/$\sqrt{\rm Hz}$). This converts into a frequency noise floor of $S_{\nu}\simeq 5\times 10^{-3}$ Hz2/Hz. When using the transimpedance amplifier, the electronic noise increases to typically $0.1$ Hz2/Hz, as shown with the gray solid line in Fig. 2(d, e). In addition to the electronic noise, we can be further limited by the photon shot noise. We evaluate it to $-110$ dBm/Hz for 200 $\mu$W of power onto the photo-detectors, leading to an equivalent frequency noise floor of $0.2~{}\rm Hz^{2}/Hz$, represented by the gray dashed line. Together, the electronic and photon shot noise add up to a detection limit represented in dark blue, giving 32 dB of signal-to-noise ratio.
In the Fourier frequency range of 0.1 to 10 MHz, we reached this noise floor. At lower frequencies, we however observed an increase of noise. This is explained by the remaining electronic noise of the Arduino DUE after the $\rm kHz$ low-pass filter and the sensitivity of optical fibers to external perturbations. We conclude that this simple setup is perfectly adequate for measuring fast phase noise on the microsecond timescale and faster, but not for longer timescale.